[ Joel is 8 year older than Karen, and Karen is twice as old as Lina. If Lina is z years old, how old is Joel in terms of z? ]
Found 2 solutions by stanbon, TimothyLamb:
Answer by stanbon(75887)

Joel is 8 year older than Karen, and Karen is twice as old as Lina. If Lina is x years old, how old is Joel in terms of x?
-------------------
Lina = x yrs old
Karen = 2x yrs old
Joel = 2x+8 yrs old
---------------
Cheers,
Stan H.
===============

Answer by TimothyLamb(4379)

---
x = y + 8
y = 2z
z = 0, 1, 2, ...
---
put the system of linear equations into standard form
---
x - y = 8
y - 2z = 0
z = 0, 1, 2, ...
---
by matrix method:
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
answer:
x= joel = 8 + 2z
y= karen = 2z
z= lina = 0, 1, 2, ...
---
by substitution:
---
x = y + 8
y = 2z
---
x = joel = 2z + 8
z = lina = 0, 1, 2, ...
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php